OpenStudy (anonymous):

calculate the mass of potassium dichromate required to prepare 100 cm^3 of a 2.00*10^-3 mol*dm^-3 solution....................................please show calculations

OpenStudy (anonymous):

please help out guys

OpenStudy (vincent-lyon.fr):

First, write down the chemical equation for potassium dichromate and work out its molar mass.

OpenStudy (anonymous):

Molarity = 2*10^-3 M = 2*10^-3 moles/liter. Molar mass of K2Cr2O7= 294.2 grams/mole. No. of moles of K2Cr2O7 in 100 ml of solution = 0.1*2*10^-3 = 2*10^-4 moles. => mass of K2Cr2O7 required is ( 2*10^-4 ) *294.2 = 0.0588 grams.